# DART-LLM: Dependency-Aware Multi-Robot Task Decomposition and Execution using Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated significant reasoning capabilities in robotic systems. However, their deployment in multi-robot systems remains fragmented and struggles to handle complex task dependencies and parallel execution. This study introduces the DART-LLM (Dependency-Aware Multi-Robot Task Decomposition and Execution using Large Language Models) system, designed to address these challenges. DART-LLM utilizes LLMs to parse natural language instructions, decomposing them into multiple subtasks with dependencies to establish complex task sequences, thereby enhancing efficient coordination and parallel execution in multi-robot systems. The system includes the QA LLM module, Breakdown Function modules, Actuation module, and a Vision-Language Model (VLM)-based object detection module, enabling task decomposition and execution from natural language instructions to robotic actions. Experimental results demonstrate that DART-LLM excels in handling long-horizon tasks and collaborative tasks with complex dependencies. Even when using smaller models like Llama 3.1 8B, the system achieves good performance, highlighting DART-LLM’s robustness in terms of model size. Please refer to the project website https://wyd0817.github.io/project-dart-llm/ for videos and code.
@@ I Introduction
Multi-robot systems have shown immense potential across various scenarios, such as complex construction sites and disaster response [10], due to their high flexibility and efficiency. The key advantage of these systems lies in their ability to perform tasks in parallel and collaborate to solve problems that are difficult for a single robot to handle independently. Zhao et al. proposed the RoCo framework, utilizing LLMs for high-level communication and low-level path planning [11]. Kannan et al. developed the SMART-LLM framework, specifically targeting task planning issues in multi-robot systems [12]. The comparison of the above-mentioned studies is shown in Table I. Table I. indicates that although existing research has significantly enhanced robots’ understanding of natural language and flexibility in task execution, driven by cutting-edge technologies like LLMs, there remain fragmented deployments of multi-robot systems. Moreover, these studies still face limitations in handling complex task dependencies and parallel execution. They utilize linear methods for task decomposition, making it challenging to effectively manage the interdependencies among subtasks generated from tasks with complex dependencies, which is particularly critical in multi-robot systems. The key advantage of multi-robot systems lies in their ability to collaboratively solve problems that are difficult for a single robot to handle independently. However, achieving this collaboration requires an effective mechanism that ensures logical sequencing of interdependent subtasks among multiple robots, while enabling parallel execution of independent tasks. To address these issues, we propose the DART-LLM (Dependency-Aware Multi-Robot Task Decomposition and Execution using Large Language Models) system. The main contributions are as follows:
@@
1. Integrated System Framework: An integrated system framework named DART-LLM, comprising multiple modules, is proposed. The framework includes the QA LLM module, Breakdown Function module, Actuation module, and a Vision-Language Model (VLM)-based object detection module. These modules perform Instruction Parsing and Task Decomposition, Parsing of Decomposed Tasks, Grounding in Embodiments, and Updating the Object Map Database to accomplish task decomposition and execution from natural language instructions to robotic actions.
2. Dependency-Aware Task Decomposition: A dependency-aware task decomposition method is proposed, which breaks down complex tasks into multiple interdependent subtasks. By accounting for dependencies, the DART-LLM allows execute dependent subtasks sequentially while running independent tasks in parallel, thereby enhancing logical capabilities and improving the efficiency of complex tasks.
@@ II Problem Formulation
This section presents the foundational mathematical concepts of the DART-LLM system, including the definitions of robot sets, associated skills, task decomposition, team formation, and navigation functions essential for task execution.
@@ II-A Robot Set and Skills
where $N$ is the number of robots. The set of skills for each robot is:
@@
where $S^{n}\subseteq\Delta$ and $\Delta$ is the set of all skills. Each robot $R^{n}$ has a skill set $S^{n}$.
II-B Team of Robots and Navigation Functions
@@
where $Q$ is the number of robots in the team. The skills of a team of robots are:
$\mathbb{S}_{A}=\{S_{A}^{1},S_{A}^{2},\ldots,S_{A}^{Q}\}$ (4)
The $m$ navigaiton function skills of a robot are defined as:
$\mathbb{S}_{F}=\{S_{F}^{1},S_{F}^{2},\ldots,S_{F}^{m}\}$ (5)
The set of all navigation functions for all robots is defined as
$T_{S}^{f}\subseteq\bigcup\mathbb{S}_{F}$ (6)
The skills needed by the team must satisfy:
$T_{S}^{k}\subseteq\bigcup\mathbb{S}_{A}$ (7)
Thus, the combined skills must cover all sub-task and function requirements:
$\bigcup\mathbb{S}_{AF}\supseteq T_{S}^{k}\cup T_{S}^{f}$ (8)
@@ II-C Task and Function Requirements
The high-level language instruction is represented as $\mathbb{I}$ and the environment is denoted by $\mathbb{E}$. The instruction $\mathbb{I}$ is decomposed using the QA LLM module into a set of sub-tasks. The details of this module will be described in subsection III-A:
@@
where $K$ is the number of sub-tasks, and $t_{j}$ denotes the temporal order of a sub-task, with $j\leq K$. Each task $T^{K}_{t_{j}}$ should be executable using the skills from the skill library within $\bigcup\mathbb{S}_{AF}$.
@@ II-D Dependency Relationships Between Tasks
represents the set of directed edges indicating dependencies between subtasks. An edge $(T^{k-1}_{t_{i}},T^{k}_{t_{j}})\in\mathbb{D}$ indicates that subtask $T^{k}_{t_{j}}$ depends on the successful completion of subtask $T^{k-1}_{t_{i}}$.
@@ III The DART-LLM Framework
III-A Instruction Parsing and Task Decomposition Use the QA LLM Module
High-level natural language instructions $\mathbb{I}$ serve as the input for DART-LLM. The input is first processed by the QA LLM for Instruction Parsing and Task Decomposition. The QA LLM parses the high-level instructions $\mathbb{I}$ while considering the environment $\mathbb{E}$ and the skills $\bigcup\mathbb{S}_{AF}$ possessed by all robots $\mathbb{R}$. The skills in $\bigcup\mathbb{S}_{AF}$ are derived from all robots $\mathbb{R}$ and any possible teams $\mathbb{A}$ they may form. It should be noted that during Task Decomposition, DART-LLM uses awareness of dependency relationships $\mathbb{D}$ to decompose the task into multiple subtasks $\mathbb{T}=\{T_{t_{1}}^{1},T_{t_{2}}^{1},\ldots,T_{t_{j}}^{K}\}$. This facilitates parallel execution of decomposed subtasks without mutual dependencies and clearly expresses dependencies between subtasks. Therefore, the prompt $\mathbb{P}$ for the QA LLM is defined as $\mathbb{P}=(\mathbb{I},\mathbb{E},\mathbb{R},\mathbb{S}_{A},\mathbb{S}_{F},\mathbb{F})$ , where $\mathbb{F}$ represents the few-shot example. The output of the QA LLM uses a structured standard format to handle the parsed commands and their dependencies. The standardized structure follows JSON syntax, as shown below:
@@ III-A Instruction Parsing and Task Decomposition Use the QA LLM Module
The structured standard format message contains the names of the breakdown functions (i.e., atomic action functions) along with their parameters, object keywords, and dependencies. These atomic action functions belong to $\mathbb{S}_{F}$ and $\mathbb{S}$. For specific atomic action functions, please refer to subsection III-C. The parameters of the atomic action functions are searched within the object map database. For details, please refer to subsection III-D.
III-B Parsing of Decomposed Tasks Uses the Breakdown Function Modules
The Breakdown Function Parser module receives the subtasks $\mathbb{T}=\{T_{t_{1}}^{1},T_{t_{2}}^{1},\ldots,T_{t_{j}}^{K}\}$ decomposed by the QA LLM Module and parses the instruction functions, parameters, and dependencies of all the skills used by the subtasks. The parsed content is then sent to the Breakdown Function Handler module. The Breakdown Function Handler module searches for the location information of entities using object keywords in the object map database. This database is updated using the VLM-based object detector module.
III-C Grounding in Embodiments Uses the Actuation Module
The Actuation module grounds in embodiments by mapping $\mathbb{T}=\{T_{t_{1}}^{1},T_{t_{2}}^{1},\ldots,T_{t_{j}}^{K}\}$ to $\bigcup\mathbb{S}_{AF}$. This is specifically achieved through predefined atomic action functions. All atomic action functions are asynchronously executed by the Actuation module. These atomic action functions are divided into two categories: navigation functions and robot-specific functions, corresponding to $\mathbb{S}_{F}$ and $\mathbb{S}$, respectively. The implementation of $\mathbb{S}_{F}$ is achieved through the ROS Navigation stack [14], while $\mathbb{S}$ is realized by creating specific action sets for each robot.
III-C1 Navigation Functions
The navigation functions are responsible for directing the movement and area access permissions of all robots or specific robots within a designated area. These functions include whether to avoid certain areas, return to the initial position, or move to a specified location. Table  II provides detailed descriptions of each navigation function.
TABLE II: Navigation Function Descriptions
Number Function Name Description
@@ III-C1 Navigation Functions
III-C2 Robot Functions
The robot-specific functions are tailored to the unique operational capabilities of different robot types, such as excavators and dump trucks. These functions handle tasks such as digging, unloading, and loading. Table  III provides detailed descriptions of each robot function. It should be noted that different types of robots have different atomic action function definitions.
TABLE III: Robot Function Descriptions
Number Function Name Description
@@ III-C2 Robot Functions
III-D Updating Object Map Database Uses the VLM-based object detector module
The DART-LLM updates the object map database in environment $\mathbb{E}$ using the VLM-based object detector module. This module first converts all images captured by Unmanned Aerial Vehicles (UAVs) into bird’s-eye views. These views are processed to detect and recognize objects in the environment $\mathbb{E}$ Selective search [15] is then used on the bird’s-eye views to generate candidate bounding boxes. Following region proposals, a CLIP-based model [16] is employed for identification, updating the object map database with the names, locations, and shapes of the objects.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated promising reasoning capabilities in robotics; however, their application in multi-robot systems remains limited, particularly in handling task dependencies. This paper introduces DART-LLM, a novel framework that employs Directed Acyclic Graphs (DAGs) to model task dependencies, enabling the decomposition of natural language instructions into well-coordinated subtasks for multi-robot execution. DART-LLM comprises four key components: a Question-Answering (QA) LLM module for dependency-aware task decomposition, a Breakdown Function module for robot assignment, an Actuation module for execution, and a Vision-Language Model (VLM)-based object detector for environmental perception, achieving end-to-end task execution. Experimental results across three task complexity levels demonstrate that DART-LLM achieves state-of-the-art performance, significantly outperforming the baseline across all evaluation metrics. Among the tested models, DeepSeek-r1-671B achieves the highest success rate, whereas Llama-3.1-8B exhibits superior response time reliability. Ablation studies further confirm that explicit dependency modeling notably enhances the performance of smaller models, facilitating efficient deployment on resource-constrained platforms. Please refer to the project website https://wyd0817.github.io/project-dart-llm/ for videos and code.
@@ I Introduction
Despite significant advancements in LLM-powered robotics research, most studies have primarily focused on single-robot systems, while applications in multi-robot systems remain at an early stage of exploration. Multi-robot systems demonstrate significant potential in complex task scenarios such as construction and disaster rescue, improving task execution efficiency and enabling collaborative problem-solving for tasks that are difficult for a single robot to accomplish independently [10]. However, multi-robot task planning still faces numerous challenges, particularly concerning the complexity of mobile robots and real-time execution. The RoCo framework proposed by Zhao et al. employs LLMs for high-level task communication and integrates multi-arm motion planning to accelerate trajectory generation. It also incorporates environmental feedback (e.g., collision detection), allowing the LLM to dynamically adjust task plans based on context. However, RoCo is designed for fixed-position robotic arm systems, making its task planning approach less adaptable to mobile robots or broader task scenarios [11]. Kannan et al.’s SMART-LLM has made progress in multi-robot task planning. However, it generates outputs using Python and does not explicitly handle dependencies among complex tasks. Due to the inherent complexity and rigidity of programming languages, practical testing revealed a high failure rate. Furthermore, SMART-LLM requires generating Python code that must be manually executed, lacking real-time LLM inference capabilities [12].
The comparison of the above-mentioned studies is shown in Table I. Table I highlights two key challenges that persist in multi-robot systems: (1) existing systems lack support for end-to-end real-time execution across multiple types of mobile robots; (2) current methods rely solely on the model’s intrinsic reasoning ability without explicitly representing task dependencies, which results in suboptimal performance, particularly for smaller models in handling complex tasks. For instance, RoCo [11] lacks support for mobile robots, being designed for fixed-position robotic arms, while SMART-LLM [12] does not support end-to-end real-time execution and fails to explicitly model task dependencies. To address these limitations, we propose DART-LLM, A Dependency-Aware Task Decomposition and Execution System for Multi-Robot Coordination, as illustrated in Fig. 1. This study makes the following contributions:
1. Dependency-Aware Task Decomposition Mechanism: A novel task decomposition mechanism that utilizes a Directed Acyclic Graph (DAG) to model subtask dependencies. This method significantly enhances system reasoning capabilities, particularly improving the ability of smaller models to handle complex task dependencies. It ensures proper task execution order while maximizing overall efficiency.
@@
2. End-to-End Real-Time Execution Framework: A real-time execution framework includes the QA LLM module, Breakdown Function module, Actuation module, and a Vision-Language Model (VLM)-based object detection module. These modules perform Instruction Parsing and Task Decomposition, Parsing of Decomposed Tasks, Grounding in Embodiments, and Updating the Object Map Database to accomplish task decomposition and execution from natural language instructions to robotic actions.
3. Benchmark Dataset for Construction Robot Evaluation: A dataset of 102 high-level natural language instructions for construction robot tasks, designed with strict execution order constraints and spanning three complexity levels.
Figure 2: Dependency-aware task decomposition in DART-LLM. Left: QA LLM decomposes high-level instruction into subtasks with explicit dependency lists ($d_{k}$) and assigned workers. Right: Construction of the Directed Acyclic Graph (DAG) based on dependency relationships, establishing the execution order where $T^{1}_{t_{1}}$ is executed first, followed by parallel execution of $T^{2}_{t_{2}}$ and $T^{3}_{t_{2}}$, and finally $T^{4}_{t_{3}}$ after its dependencies are satisfied.
@@ II Problem Formulation
This section presents the foundational mathematical concepts of the DART-LLM system. We formalize the definitions of robot sets and their associated skills, team formation with collaborative skill enhancement, task decomposition into sub-tasks with temporal and dependency relationships, and the skill library essential for task execution.
@@ II-A Robot Set and Skills
where $N$ is the number of robots. The set of skills for all robots are:
@@
II-B Team of Robots and Team Skills
@@
where $Q$ denotes the number of robots in the team. The skills of a team of robots are expressed as:
$\mathbb{S}_{A}=\{S_{A}^{1},S_{A}^{2},\ldots,S_{A}^{\Omega}\}$ (4)
where $\Omega\geq Q$. Here, $\Omega$ represents the total number of skills, including both the individual skills of each robot and the new skills generated through robot collaboration. The set of all skills across all teams formed by $N$ robots is:
$\Delta=\bigcup\mathbb{S}_{A}$ (5)
It should be noted that $\Delta$ includes $\mathbb{S}$, the set of all individual robot skills.
@@ II-C Task and Function Requirements
The high-level language instruction is represented as $\mathbb{I}$, and the environment is denoted by $\mathbb{E}$. The instruction $\mathbb{I}$ is decomposed using the QA LLM module into a set of sub-tasks. A set of sub-tasks is defined as:
@@
where $K$ is the number of sub-tasks, and $t_{j}$ denotes the temporal order of a sub-task. Since sub-tasks may be executed simultaneously, we have $j\leq K$. Each task $T^{k}_{t_{j}}$ should be executable using the skills from the skill library within $\Delta$.
@@ II-D Dependency Relationships Between Tasks
represents the set of directed edges indicating dependencies between subtasks. An edge $\mathbb{D}_{k}=(T^{k-1}_{t_{i}},T^{k}_{t_{j}})\in\mathbb{D}$ indicates that subtask $T^{k}_{t_{j}}$ depends on the successful completion of subtask $T^{k-1}_{t_{i}}$.
@@ III The DART-LLM Framework
To illustrate the workflow of DART-LLM, Algorithm 1 presents the pseudocode for the system’s dependency-aware multi-robot task decomposition and execution process. This algorithm outlines the system’s processing of natural language instructions through multiple stages, including instruction parsing, dependency-aware multi-robot task decomposition, dependency-aware task parsing and robot assignment, and task execution with actuation. Following this high-level overview, we detail each module shown in Fig. 1 and explain how they implement the different stages of Algorithm 1 in the subsequent subsections.
Algorithm 1 Dependency-Aware Multi-Robot Task Decomposition and Execution
0:  Natural language instruction $\mathbb{I}$, environment information $\mathbb{E}$, robot set $\mathbb{R}$ with skill sets $\mathbb{S}$ and a few-shot example set $\mathbb{F}$
1:  Instruction Parsing
2:  Provide QA LLM module with $\mathbb{I}$ for parsing
3:  Generate prompt $\mathbb{P}=(\mathbb{I},\mathbb{E},\mathbb{R},\mathbb{S},\mathbb{F})$ with explicit instruction that robot skills can combine through teamwork
4:  Obtain parsed command structure in JSON format, including subtasks, robot assignments, dependencies, and object keywords
5:  Dependency-Aware Multi-Robot Task Decomposition
6:  Decompose $\mathbb{I}$ into subtasks $\mathbb{T}=\{T^{1}_{t_{1}},T^{1}_{t_{2}},\ldots,T^{K}_{t_{j}}\}$ with QA LLM, where each $T^{k}_{t_{j}}$ corresponds to an atomic skill in $\mathbb{S}$
7:  Define dependency relationships $\mathbb{D}\subseteq\mathbb{T}\times\mathbb{T}$ during decomposition
8:  Represent dependencies using Directed Acyclic Graph $G=(\mathbb{T},\mathbb{D})$, where $\mathbb{D}$ enables task coordination and implicit team formation
9:  Dependency-Aware Task Parsing and Robot Assignment
10:  Initialize object map database with object locations and attributes using VLM-based object detector
11:  for each subtask $T^{k}_{t_{j}}$ in topological order of DAG $G$ do
12:     if $T^{k}_{t_{j}}$ has no dependencies in $\mathbb{D}$ then
13:        Execute $T^{k}_{t_{j}}$ in parallel
14:     else
15:        Wait for all dependencies of $T^{k}_{t_{j}}$ in $\mathbb{D}$ to complete
16:        Execute $T^{k}_{t_{j}}$
17:     end if
18:  end for
19:  for each subtask $T^{k}_{t_{j}}\in\mathbb{T}$ do
20:     if QA LLM specified a particular robot for $T^{k}_{t_{j}}$ then
21:        Assign $T^{k}_{t_{j}}$ to the specified robot from $\mathbb{R}$ with corresponding skill in $\mathbb{S}$
22:     else
23:        Use Breakdown function parser to find available robot of specified type from $\mathbb{R}$ with matching skill
24:     end if
25:  end for
26:  Task Execution using Actuation Module
27:  for each parsed subtask $T^{k}_{t_{j}}\in\mathbb{T}$ do
28:     Execute assigned robot-specific atomic skill based on $\mathbb{S}$
29:  end for
30:  Update Object Map Database
31:  Continuously update object map based on real-time VLM object detector feedback
@@
III-A Instruction Parsing and Dependency-Aware Multi-Robot Task Decomposition Using the QA LLM Module
This subsection details steps 1-8 of Algorithm 1, focusing on how high-level natural language instructions $\mathbb{I}$ are processed by the QA LLM module and decomposed into executable subtasks with dependency relationships.
High-level natural language instructions $\mathbb{I}$ serve as the input for DART-LLM. The input is first processed by the QA LLM for Instruction Parsing and Dependency-Aware Task Decomposition. The QA LLM parses the high-level instructions $\mathbb{I}$ while considering the environment $\mathbb{E}$ and the skills $\mathbb{S}$ possessed by all robots $\mathbb{R}$.
In this module, we employ a DAG-based task decomposition system that decomposes a given instruction $\mathbb{I}$ into subtasks while also generating a dependency list for each subtask to ensure correct execution sequencing. As shown in Fig. 2, when instruction $\mathbb{I}$ is decomposed into a set of subtasks $T$ by QA LLM, each subtask $T^{k}_{t_{j}}$ is assigned a dependency list $d_{k}$, which specifies all prerequisite subtasks that must be completed before executing $T^{k}_{t_{j}}$. The system then determines the execution order $t_{i}$ automatically based on these dependencies and constructs a directed edge set $D_{k}$ for each $d_{k}$, thereby forming a subtask execution graph as illustrated. Finally, the system executes subtasks strictly in the specified order until the overall task objective is achieved.
During Task Decomposition, DART-LLM uses awareness of dependency relationships $\mathbb{D}$ to decompose the task into multiple subtasks $\mathbb{T}=\{T_{t_{1}}^{1},T_{t_{2}}^{1},\ldots,T_{t_{j}}^{K}\}$, where each subtask corresponds to an atomic skill in $\mathbb{S}$. This facilitates parallel execution of decomposed subtasks without mutual dependencies and clearly expresses dependencies between subtasks.
Therefore, the prompt $\mathbb{P}$ for the QA LLM is defined as $\mathbb{P}=(\mathbb{I},\mathbb{E},\mathbb{R},\mathbb{S},\mathbb{F})$ , where $\mathbb{F}$ represents the few-shot example, with explicit instructions that robot skills can combine through teamwork.
The output of the QA LLM uses a structured standard format to handle the parsed commands and their dependencies. The standardized structure follows JSON syntax, as shown below:
@@ III-A Instruction Parsing and Dependency-Aware Multi-Robot Task Decomposition Using the QA LLM Module
The structured standard format message contains the names of the breakdown functions (i.e., atomic action skills) along with their parameters, object keywords, and dependencies. These atomic action skills belong to $\mathbb{S}$. For specific atomic action skills, please refer to subsection III-C. The parameters of the atomic action skills are searched within the object map database. For details, please refer to subsection III-D.
III-B Dependency-Aware Task Parsing and Robot Assignment Using the Breakdown Function Modules
This subsection elaborates on the implementation details of steps 9-25 of Algorithm 1, particularly how the system processes tasks according to their dependencies and assigns robots to subtasks.
The system first initializes the object map database with object locations and attributes using the VLM-based object detector. This provides the environmental context necessary for task execution and robot coordination.
Following the DAG structure established during task decomposition, the system processes each subtask $T^{k}_{t_{j}}$ in topological order of the graph $G$. This ordering ensures that subtasks are executed according to their dependencies: - Subtasks without dependencies are executed in parallel, maximizing efficiency - Subtasks with dependencies wait for their prerequisite tasks to complete before execution
The Breakdown Function Parser module then processes each subtask $T^{k}_{t_{j}}\in\mathbb{T}$ for robot assignment. If the QA LLM has specified a particular robot for a subtask, that robot is directly assigned to the task. Otherwise, the Breakdown Function Parser finds an available robot of the specified type from $\mathbb{R}$ with the matching skill required for the subtask.
This dependency-aware approach to task parsing and robot assignment ensures that complex tasks requiring collaboration between multiple robots are executed efficiently while respecting the necessary order of operations defined by the dependency relationships in $\mathbb{D}$.
III-C Task Execution and Grounding Using the Actuation Module
This subsection corresponds to steps 26-29 of Algorithm 1, following the dependency-aware execution and robot assignment described previously, detailing how the system grounds instructions in robot actions through the actuation module.
The Actuation module grounds the instructions in embodiments by executing each parsed subtask $T^{k}_{t_{j}}\in\mathbb{T}$ using the corresponding atomic skill from $\mathbb{S}$. All atomic action skills are asynchronously executed by the Actuation module.
These atomic action skills are divided into two categories: navigation skills and robot-specific skills, both contained within $\mathbb{S}$. The implementation of navigation skills is achieved through the ROS Navigation stack [14], while robot-specific skills are realized by creating specific action sets for each robot.
III-C1 Navigation Skills
The navigation skills are responsible for directing the movement and area access permissions of all robots or specific robots within a designated area. These skills include whether to avoid certain areas, return to the initial position, or move to a specified location. Table  II provides detailed descriptions of each navigation skill.
TABLE II: Navigation Skill Descriptions
Number Skill Name Description
@@ III-C1 Navigation Skills
III-C2 Robot Skills
The robot-specific skills are tailored to the unique operational capabilities of different robot types, such as excavators and dump trucks. These skills handle tasks such as digging, unloading, and loading. Table  III provides detailed descriptions of each robot skill. It should be noted that different types of robots have different atomic action skill definitions.
TABLE III: Robot Skill Descriptions
Number Skill Name Description
@@ III-C2 Robot Skills
III-D Update Object Map Database Using the VLM-based Object Detector Module
This subsection explains steps 30-31 of Algorithm 1, describing how the system continuously updates its object map database through real-time sensing, which is crucial for maintaining an accurate representation of the environment throughout task execution.
The DART-LLM updates the object map database in environment $\mathbb{E}$ using the VLM-based object detector module. The object detector focuses on identifying the object keywords that were extracted during QA LLM’s instruction parsing phase, ensuring that all entities relevant to the current task are properly detected and tracked.
This module first converts all images captured by Unmanned Aerial Vehicles (UAVs) into bird’s-eye views. These views are processed to detect and recognize objects in the environment $\mathbb{E}$ Selective search [15] is then used on the bird’s-eye views to generate candidate bounding boxes. Following region proposals, a CLIP-based model [16] is employed for identification, updating the object map database with the names, locations, and shapes of the objects. The updated information is then used by the Breakdown Function Handler to ensure accurate parameter resolution for the atomic skills during execution.